i need a answer to:
3x-y=2
x+2y=10

Answer by vidhyak(98)


3x -y =2 ......eq 1
x +2y =10 ......eq 2

Multiply eq1 by 2
6x -2y =4
x +2y =10

Add the above equations
7x =14
x=2

Substitute x = 2 in eq 1
3x -y =2
3(2) -y =2
6 - y =2
-y = -4
y = 4

x = 2, y = 4